# Making Differential Privacy Easier to Use for Data Controllers using a Privacy Risk Indicator

### v4

@@ Abstract.
PVLDB Reference Format:
PVLDB, 14(1): XXX-XXX, 2020.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at %leave␣empty␣if␣no␣availability␣url␣should␣be␣sethttps://github.com/youny626/DP_test.
@@ 1. Introduction
In this paper, we provide controllers with additional information that complements the worst-case measure of disclosure risk given by $\epsilon$ by introducing a relative, within-dataset measure of disclosure risk. We call this Relative Disclosure Risk Indicator (RDR). Given the controller’s dataset and an analyst’s query, the RDR indicates the expected disclosure risk of a specific within-dataset individual relative to other within-dataset individuals upon releasing a DP output; this is different from another measure of disclosure risk defined by Lee and Clifton (Lee and Clifton, 2011), which is the maximum probability of identifying any within-dataset individuals. With RDR, controllers understand the relative disclosure risk of any two within-dataset individuals, e.g., by comparing the most at-risk individual (highest RDR) and the least at-risk individual (lowest RDR). RDR communicates to controllers the impact of choosing $\epsilon$ on the within-dataset individuals, and we hypothesize that this additional information helps controllers make more informed choices of $\epsilon$. To test our hypothesis, we run a between-subjects user study where we ask participants to choose $\epsilon$ for a given query and dataset, and we present the treatment group with the additional information of the within-dataset individuals’ RDRs. We observe that the treatment group makes more consistent choices of $\epsilon$ compared to the control group, and all participants in the treatment group used RDRs in their decision-making process, thus demonstrating that RDR is effective in helping controllers choose $\epsilon$.
@@
Equipped with RDRs, we make several technical contributions. First, we develop the key insight that controllers can express their privacy preferences on the within-dataset individuals as a function of the within-dataset individuals’ RDRs. For instance, controllers can specify the maximum permissible difference between the most at-risk and least at-risk within-dataset individuals. Second, we design an algorithm, called Find-$\epsilon$-from-RDR, that takes controllers’ privacy preferences in terms of the within-dataset individuals’ RDRs and derives an $\epsilon$ that satisfies those preferences. Because RDRs are data-dependent, they are only visible to controllers and never to analysts, and the derived $\epsilon$ cannot be released to the analysts either. Thus third, we introduce an alternative algorithm, called Find-and-release-$\epsilon$-from-RDR, to enable DP release of the derived $\epsilon$. Lastly, in practice, analysts may submit multiple queries to the controller, and the controller needs to set a total privacy budget of their dataset. Choosing this budget suffers from the same difficulty as choosing $\epsilon$ (Dwork et al., 2019; Cummings et al., 2021; Garrido et al., 2022). We propose an extension to both algorithms that allows controllers to dynamically determine whether to allow or deny answering analysts’ queries, hence circumventing the need to choose a total budget. We construct a privacy odometer (Rogers et al., 2016) to provide an end-to-end DP guarantee for our solution.
@@ 3.1. Why Choosing $\epsilon$ is Hard and RDR Overview
When controllers decide which $\epsilon$ to choose for an analyst’s query, they may naturally ask the question: how would choosing this $\epsilon$ affect the privacy of the within-dataset individuals I am trying to protect? Unfortunately, the value of $\epsilon$ by itself does not provide a straightforward answer. Because $\epsilon$ quantifies a data-independent worst-case privacy guarantee, it is challenging to translate this privacy guarantee to a practical measure of disclosure risk of specific within-dataset individuals. The global and probabilistic nature of DP gives a robust privacy guarantee, but it also makes it difficult for controllers to reason about the concrete privacy implications for each within-dataset individual when they apply DP in practice.
We hypothesize that controllers can only comfortably choose $\epsilon$ if they understand how each within-dataset individual’s privacy is affected by that choice. Therefore, we design an indicator that communicates to controllers the relative disclosure risk of each within-dataset individual as an effect of choosing a specific $\epsilon$; we call it the Relative Disclosure Risk Indicator (RDR). Comparing two within-dataset individuals’ RDRs, the one with a higher RDR has greater disclosure risk than the other (upon releasing the DP output produced by the chosen $\epsilon$). We design RDR to be a relative indicator so that controllers can easily compare the values and assess the significance of differences; studies across various fields have shown that relative measures offer enhanced interpretability and practical advantages over absolute measures (Keiningham et al., 2015; Citrome, 2010; Reich et al., 2016). The RDR complements any information controllers already have access to, and it remains private information to controllers only.
@@ 3.2. Output-dependent RDR
Given query $q:\mathcal{X}^{n}\rightarrow\mathbb{R}^{k}$, dataset $x\in\mathcal{X}^{n}$, $\epsilon>0$, the Laplace Mechanism is defined as:
@@
$RDR_{i}^{(o)}>RDR_{j}^{(o)}$ means $\lVert o-q(x_{-i})\rVert_{1}$ $\lVert o-q(x_{-j})\rVert_{1}$ (the oRDR is in $\ell_{1}$ norm since Laplace Mechanism is used), thus
@@
Given query $q:\mathcal{X}^{n}\rightarrow\mathbb{R}^{k}$, dataset $x\in\mathcal{X}^{n}$, $\epsilon,\delta>0$, the Gaussian Mechanism is defined as:
@@
where $Z_{i}$ are i.i.d random variables drawn from the Gaussian distribution $\mathcal{N}(\sigma^{2})$ with center 00 and variance $\sigma^{2}=\frac{2\Delta_{2}^{2}\ln(1.25/\delta)}{\epsilon^{2}}$  [^4] The global sensitivity of q in the $\ell_{2}$ norm is defined as $\Delta_{2}=\max\lVert q(x)-q(x^{\prime})\rVert_{2}$, where $x$ and $x^{\prime}$ are neighboring datasets in $\mathcal{X}^{n}$.. The Gaussian Mechanism satisfies $(\epsilon,\delta)$-DP.
@@
$RDR_{i}^{(o)}>RDR_{j}^{(o)}$ means $\lVert o-q(x_{-i})\rVert_{2}$ $\lVert o-q(x_{-j})\rVert_{2}$; the oRDR is in $\ell_{2}$ norm under Gaussian Mechanism.
@@ 3.3. RDR: Formal Definition
$RDR_{i}=\mathbb{E}_{o\sim\mathcal{M}(x,q,\epsilon,\delta)}[PRI_{i}^{(o)}].$
@@
The $RDR_{i}$ under Laplace Mechanism is
@@
The $RDR_{i}$ under Gaussian Mechanism is
@@
$k$ is the output dimension of query $q$. As noted in Definition 3.3 and 3.4, $\Delta_{1}/\epsilon$ is the noise scale of the Laplace distribution, and $\sigma^{2}$ is the variance of the Gaussian distribution. We omit details of the calculation due to space constraints and include them in the associated technical report.
@@ 4. Deriving Epsilon based on RDR
In this section, we describe the intuition of using RDR to express controllers’ privacy preferences over the within-dataset individuals (Section 4.1), then introduce an algorithm to choose $\epsilon$ by using RDR as a more intuitive proxy (Section 4.2). Lastly, we discuss the algorithm’s complexity and extensions (Section 4.3).
@@ 4.1. Using RDR to Express Privacy Preferences
The RDR complements the worst-case bound of disclosure risk given by $\epsilon$ and becomes an additional indicator of relative disclosure risk associated with each within-dataset individual. A controller may express their privacy preferences over the within-dataset individuals in terms of their RDRs, and choose an $\epsilon$ such that the corresponding RDRs satisfy their privacy preferences. Next, we convey this intuition through an example.
@@
The controller chooses $\epsilon$ based on their privacy preferences in terms of the ratio between the maximum and maximum RDR of the within-dataset patients. Between $\epsilon=1$ and $\epsilon=0.1$, the controller might be more comfortable choosing $\epsilon=0.1$ since the within-dataset patients’ RDRs are closer together (the ratio between the minimum and maximum RDR increases to $\frac{10}{11}$ from $\frac{1}{2}$). Between $\epsilon=0.1$ and $\epsilon=0.01$, although the latter produces even closer RDRs (the ratio becomes $\frac{100}{101}$), it comes at the cost of worse output accuracy. Thus one sensible choice is to choose $\epsilon=0.1$ as it yields higher accuracy for the analyst receiving the output.
Key idea. When a controller computes a query by applying a DP mechanism with a specific $\epsilon$, observing the RDRs associated with the within-dataset individuals communicates to the controller the relative disclosure risk of each within-dataset individual if the output was released. The controller may then express their privacy preferences over the within-dataset individuals by specifying a function of the individuals’ RDRs, e.g., the ratio between the minimum RDR ($RDR_{min}$) and maximum RDR ($RDR_{max}$), and a threshold over the function’s output, e.g., a percentage $0\leq\tau_{p}\leq 1$. The controller’s privacy preferences are satisfied when $RDR_{min}/RDR_{max}\geq\tau_{p}$. In the above example, if the controller sets $\tau_{p}=0.9$, their privacy preferences are satisfied when $\epsilon=0.1$ since $\frac{10}{11}\geq 0.9$.
@@ 4.2. The Find-$\epsilon$-from-RDR Algorithm
We design the Find-$\epsilon$-from-RDR algorithm based on the intuition we presented in the above example. By letting the controller express their privacy preferences as a function of within-dataset individuals’ RDRs, we can compute the within-dataset individuals’ RDRs for different candidate $\epsilon$ values and find a suitable $\epsilon$ that satisfies the controller’s privacy preferences. The Find-$\epsilon$-from-RDR algorithm represents the controller’s privacy preferences in terms of the function $RDR_{min}/RDR_{max}$ and asks the controller to specify the threshold $\tau_{p}$. It then finds the largest $\epsilon$ such that $RDR_{min}/RDR_{max}\geq\tau_{p}$, because if controllers’ privacy preferences can be satisfied, larger $\epsilon$ is preferred as it leads to more accurate output.
Unlike $\epsilon$, the parameter $\tau_{p}$ is intuitive for controllers to interpret and directly conveys how close the individuals’ disclosure risk should be. For instance, setting $\tau_{p}=1$ means the controller wants all within-dataset individuals to have the same disclosure risk. Furthermore, controllers may also use other functions to express their privacy preferences, such as setting a threshold for the maximum variance among RDRs, as we will show in Section 5.1.
Input : DP mechanism $\mathcal{M}$, dataset $x$, query $q$, threshold $\tau_{p}$, set of candidate $\epsilon$ values $\mathcal{E}$, failure parameter $\delta$
@@
1
[…]
3
[…]
5
6for $x_{i}\in x^{\prime}$ do
7       if $x_{i}\notin PIS$ then
8             $PIS[x_{i}]\leftarrow\textsc{compute-per-instance-sensitivity}(x^{\prime},x_{i},q,M)$
9
10
11/* Controllers may observe the computed per-instance sensitivity of each individual first and set $\tau_{p}$ based on their privacy preferences */
12
13for $\epsilon\in\mathcal{E}$ do
14
15      $RDR\leftarrow list()$
16       for $x_{i}\in x^{\prime}$ do
17             $RDR_{i}\leftarrow\textsc{compute-rdr}(PIS[x_{i}],M,q,\epsilon,\delta)$
18             $RDR.add(RDR_{i})$
19
20      $RDR_{min},RDR_{max}\leftarrow\min(RDR),\max(RDR)$
21
22      /* Check whether RDR satisfy controller’s privacy preference */
23       if $RDR_{min}/RDR_{max}\geq\tau_{p}$ then
24
25            /* Found suitable $\epsilon$. Compute DP output using chosen $\epsilon$/
26             $o\leftarrow\mathcal{M}(x^{\prime},q,\epsilon,\delta)$
27             return $o$
28
@@ 26             $o\leftarrow\mathcal{M}(x^{\prime},q,\epsilon,\delta)$
Detailed Algorithm. Algorithm 1 shows the pseudo-code for finding $\epsilon$ from RDR and releasing DP output $o$. The controller will provide the set of candidate $\epsilon$ to choose from; for example, the values could range from 10 to 0.01 in a fixed interval. The algorithm first projects relevant query attributes from the dataset (line 1) to avoid unnecessary recomputation of duplicate data (e.g., if the query only involves the age attribute, the algorithm only needs to compute the RDRs of hundreds of unique ages, regardless of the original size of the dataset). Then, for each unique within-dataset individual’s data $x_{i}$, the algorithm computes its per-instance sensitivity, which depends on the query and which DP mechanism will be used (line 1 - 1); the per-instance sensitivity will be used to compute the RDR of each within-dataset individual.
At this point, the controller may observe each within-dataset individual’s per-instance sensitivity and set an appropriate $\tau_{p}$ based on their privacy preferences over the within-dataset individuals.
Next, the algorithm iterates over the set of candidate $\epsilon$ values in descending order (line 1). For each $\epsilon$, the algorithm computes the RDR of each individual (line 1 - 1), and $RDR_{min},RDR_{max}$ (line 1). If $RDR_{min}/RDR_{max}\geq\tau_{p}$, then the algorithm has found a suitable $\epsilon$ that satisfies the controller’s privacy preferences, so it computes a DP output using the chosen $\epsilon$ and returns the output (line 1 - 1). Note that the algorithm does not release the chosen $\epsilon$, since $\epsilon$ is chosen based on data-dependent RDRs and only the controller can access the RDRs and chosen $\epsilon$.
@@
Dataflows and leakage. It is critically important to understand what information is released to whom. Figure 3 shows the involved dataflows in a scenario where a controller uses the Find-$\epsilon$-from-RDR algorithm to find $\epsilon$ and release query output to the analyst. Individuals send their data to the controller; this dataflow remains unchanged from the standard central DP deployment shown in Figure 1. The analyst submits a query to the controller. The controller expresses their privacy preferences in terms of $\tau_{p}$ (and not in terms of $\epsilon$) and runs the algorithm to release an output produced by a DP mechanism (Laplace Mechanism or Gaussian Mechanism in our work) with an $\epsilon$ selected by the algorithm. At this point, no information has been returned to the analyst yet. The RDRs of within-dataset individuals are never released to the analyst. The DP output is only released with the controller’s approval. The chosen $\epsilon$ is also not released since $\epsilon$ is chosen based on data-dependent RDRs, and releasing it directly jeopardizes the DP guarantee. In Section 5, we explain a modification of the algorithm that permits finding and releasing $\epsilon$ while satisfying DP.
@@
Claim: The Find-$\epsilon$-from-RDR algorithm satisfies $(\epsilon,\delta)$-DP. The algorithm only releases the output $o$ computed using an $(\epsilon,\delta)$-DP mechanism ($\delta=0$ when $\mathcal{M}$ is Laplace Mechanism), where $\epsilon$ is found by the algorithm that satisfies the controller’s privacy preferences. If no suitable $\epsilon$ is found, the algorithm does not release anything; in other words, $\epsilon=0$. Therefore, the algorithm satisfies $(\epsilon,\delta)$-DP.
4.3. Discussion and Extensions
Algorithm complexity and optimizations. The algorithm has a worst-case time complexity of $O(n\lvert\mathcal{E}\rvert)$, where $n$ is the number of individual records in the dataset. Computing the RDRs requires computing the per-instance sensitivity first. When the number of individual records is high, this becomes a performance bottleneck: using memoization saves computation if the number of unique projected records is smaller than $n$. In addition, the computation of each individual’s per-instance sensitivity can be parallelized since they are independent of one another, i.e., the problem is embarrassingly parallel. Thus, we implement parallelism and compute the per-instance sensitivity for a chunk of records in each process. We show in Section 6.2 that the algorithm is both efficient and scalable.
Expressing privacy preferences. By default, the algorithm considers the RDRs of all within-dataset individuals when finding $\epsilon$. A modification of our algorithm permits the controller to specify their privacy preferences in different granularity by selecting subsets of within-dataset individuals. For example, a controller can divide the dataset $x$ consisting of individuals’ ages to the subset of individuals over 60 ($x_{age>60}$) and those younger ($x_{age\leq 60}$), and specify that the two groups should have similar disclosure risk. This preference can easily be incorporated into the algorithm: instead of comparing the global $RDR_{min}$ and $RDR_{min}$ of all within-dataset individual, the controller can compare the median RDR of each group, $RDR_{med}^{age>60}$ and $RDR_{med}^{age\leq 60}$, and choose an $\epsilon$ such that $RDR_{med}^{age>60}/RDR_{med}^{age\leq 60}\geq\tau_{p}$.
Communicating Privacy-Accuracy Trade-off. In some circumstances, the controller may want to consider the accuracy of the released output when deciding on the level of privacy protection they want to offer each within-dataset individual. Prior work such as DPComp (Hay et al., 2016) and ViP (Nanayakkara et al., 2022) have explored different ways to visualize the privacy-accuracy trade-off in DP. The RDR complements these existing interfaces by providing more information on the relationship between output accuracy and within-dataset individual’s disclosure risk.
@@ 4.3. Discussion and Extensions
5. Releasing Epsilon Privately and Bounding Privacy Loss

### v5

# Within-Dataset Disclosure Risk for Differential Privacy

@@ Abstract.
Differential privacy, data controller, privacy loss budget
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Security and privacy Privacy protections
@@ 1. Introduction
In this paper, we provide controllers additional information that complements the worst-case measure of disclosure risk by introducing a relative, within-dataset measure of disclosure risk. We call this Relative Disclosure Risk Indicator (RDR). Given the controller’s dataset and analyst’s query, the RDR indicates the expected disclosure risk of a specific within-dataset individual relative to other within-dataset individuals upon releasing a DP output; this is different from another absolute measure of disclosure risk defined by Lee and Clifton (Lee and Clifton, 2011), which is the maximum probability of identifying any within-dataset individuals. With RDR, controllers understand the impact of choosing $\epsilon$ by observing the distribution of disclosure risks among the within-dataset individuals. We hypothesize that RDR helps controllers make more informed choices of $\epsilon$ as this additional information empowers controllers to make explicit decisions on their preferred risk distribution over the within-dataset individuals. To test our hypothesis, we run a between-subjects user study where we ask participants to choose $\epsilon$ for a given query and dataset, and we present the treatment group with the additional information of the within-dataset individuals’ RDRs. We observe that the treatment group makes more consistent choices of $\epsilon$ compared to the control group, and all participants in the treatment group used RDRs in their decision-making process, thus demonstrating that RDR is effective in helping participants choose $\epsilon$.
@@
Equipped with RDRs, we make several technical contributions. First, we develop the key insight that controllers can express their privacy preferences on the within-dataset individuals as a function of the within-dataset individuals’ RDRs and a threshold over the function output. For instance, controllers can specify the maximum permissible difference between the most at-risk and least at-risk within-dataset individuals. Second, we design an algorithm, called Find-$\epsilon$-from-RDR, that takes controllers’ privacy preferences in terms of the within-dataset individuals’ RDRs and derives an $\epsilon$ that satisfies those preferences. Because RDRs are data-dependent, they are only visible to controllers and never to analysts, and the derived $\epsilon$ cannot be released to the analysts either. Thus third, we introduce an alternative algorithm, called Find-and-release-$\epsilon$-from-RDR, to enable DP release of the derived $\epsilon$. Lastly, in practice, analysts may submit multiple queries to the controller, and the controller needs to set a total privacy budget of their dataset. Choosing this budget suffers from the same difficulty as choosing $\epsilon$ (Dwork et al., 2019; Cummings et al., 2021; Garrido et al., 2022). We propose an extension to both algorithms that allows controllers to dynamically determine whether to allow or deny answering analysts’ queries, hence circumventing the need to choose a total budget. We construct a privacy odometer (Rogers et al., 2016) to provide an end-to-end DP guarantee for our solution.
@@ 3.1. Why Choosing $\epsilon$ is Hard and RDR Overview
When controllers decide which $\epsilon$ to choose for an analyst’s query, they may naturally ask the question: how would choosing this $\epsilon$ affect the privacy of the within-dataset individuals I am trying to protect? However, the value of $\epsilon$ itself does not provide a straightforward answer. Because $\epsilon$ quantifies a data-independent worst-case privacy guarantee, it is challenging to translate this privacy guarantee to a practical measure of disclosure risk of specific within-dataset individuals. The global and probabilistic nature of DP gives a robust privacy guarantee, but it also makes it difficult for controllers to reason about the concrete privacy implications for each within-dataset individual when they apply DP in practice.
We hypothesize that controllers can only comfortably choose $\epsilon$ if they understand how each within-dataset individual’s privacy is affected by that choice. A fundamental challenge in deploying differential privacy is translating the abstract guarantee of $\epsilon$ into a concrete measure of disclosure risk. Lee and Clifton address this challenge by defining disclosure risk as the maximum posterior probability of re-identifying any within-dataset individuals (Lee and Clifton, 2011). This provides a single, absolute measure of the worst-case disclosure risk, establishing a ceiling on the harm to the single most vulnerable within-dataset individual. While valuable, this single-point estimate conceals the underlying distribution of risk below that ceiling. A controller informed only of the maximum disclosure risk of the within-dataset individuals cannot distinguish between the scenario where the risk is borne equally by all individuals and one where it is dangerously concentrated on a vulnerable few. Therefore, our goal is to design a relative, per-individual indicator that complements this absolute, worst-case measure of disclosure risk; we call it the Relative Disclosure Risk Indicator (RDR). Crucially, the RDR gives controllers a view into the risk distribution of the within-dataset individuals. For instance, by comparing the RDRs of the most and least at-risk individuals, or by examining the overall variance of the RDR distribution, a controller can make a more informed judgment about whether a given $\epsilon$ provides an equitable level of protection across the within-dataset individuals. The RDR complements any information controllers already have access to, and it remains private information to controllers only.
@@ 3.2. Output-dependent RDR
Given query $q:\mathcal{X}^{n}\rightarrow\mathbb{R}^{k}$, dataset $x\in\mathcal{X}^{n}$, $\epsilon>0$, Laplace Mechanism is defined as:
@@
$RDR_{i}^{(o)}>RDR_{j}^{(o)}$ means $\lVert o-q(x_{-i})\rVert_{1}$ ¿ $\lVert o-q(x_{-j})\rVert_{1}$ (the oRDR is in $\ell_{1}$ norm since Laplace Mechanism is used), thus
@@
Given query $q:\mathcal{X}^{n}\rightarrow\mathbb{R}^{k}$, dataset $x\in\mathcal{X}^{n}$, $\epsilon,\delta>0$, Gaussian Mechanism is defined as:
@@
where $Z_{i}$ are i.i.d random variables drawn from the Gaussian distribution $\mathcal{N}(\sigma^{2})$ with center 0 and variance $\sigma^{2}=\frac{2\Delta_{2}^{2}\ln(1.25/\delta)}{\epsilon^{2}}$  [^4] The global sensitivity of q in the $\ell_{2}$ norm is defined as $\Delta_{2}=\max\lVert q(x)-q(x^{\prime})\rVert_{2}$, where $x$ and $x^{\prime}$ are neighboring datasets in $\mathcal{X}^{n}$.. The Gaussian Mechanism satisfies $(\epsilon,\delta)$-DP.
@@
$RDR_{i}^{(o)}>RDR_{j}^{(o)}$ means $\lVert o-q(x_{-i})\rVert_{2}$ ¿ $\lVert o-q(x_{-j})\rVert_{2}$; the oRDR is in $\ell_{2}$ norm under Gaussian Mechanism.
@@ 3.3. RDR: Formal Definition
$RDR_{i}=\mathbb{E}_{o\sim\mathcal{M}(x,q,\epsilon,\delta)}[RDR_{i}^{(o)}].$
@@
$RDR_{i}$ upper bound under Laplace Mechanism is
@@
$RDR_{i}$ upper bound under Gaussian Mechanism is
@@
$k$ is the output dimension of query $q$. As noted in Definition 3.3 and 3.4, $\Delta_{1}/\epsilon$ is the noise scale of Laplace distribution, and $\sigma^{2}$ is the variance of Gaussian distribution. We omit details of the calculation due to space constraints and include them in the Appendix.
@@ 4. Deriving Epsilon based on RDR
In this section, we describe the key idea of using RDR to express controllers’ privacy preferences over the within-dataset individuals (Section 4.1), then introduce an algorithm to choose $\epsilon$ by using RDR as a more intuitive proxy (Section 4.2).
@@ 4.1. Using RDR to Express Privacy Preferences
The RDR complements the worst-case bound of disclosure risk given by $\epsilon$ and becomes an additional indicator of relative disclosure risk associated with each within-dataset individual. A controller may express their privacy preference over the within-dataset individuals in terms of their RDRs, and choose an $\epsilon$ such that the corresponding RDRs satisfy their privacy preference. Next, we convey this intuition through an example.
@@
The controller chooses $\epsilon$ based on their privacy preference in terms of the ratio between the maximum and maximum RDR of the within-dataset patients. Between $\epsilon=1$ and $\epsilon=0.1$, the controller might be more comfortable choosing $\epsilon=0.1$ since the within-dataset patients’ RDRs are closer together (the ratio between the minimum and maximum RDR increases to $\frac{10}{11}$ from $\frac{1}{2}$). Between $\epsilon=0.1$ and $\epsilon=0.01$, although the latter produces even closer RDRs (the ratio becomes $\frac{100}{101}$), it comes at the cost of worse output accuracy. Thus one sensible choice is to choose $\epsilon=0.1$ as it yields higher accuracy for the analyst receiving the output.
Key idea. When a controller computes a query by applying a DP mechanism with a specific $\epsilon$, observing the RDRs associated with the within-dataset individuals reveals the distribution of the individuals’ disclosure risks. This enables the controller to formally express their privacy preference over the within-dataset individuals by specifying a preference function $f:\mathbb{R}^{k}\rightarrow\mathbb{R}$ in terms of individuals’ RDRs, e.g., $f(RDR)=RDR_{min}/RDR_{max}$, and a threshold $\tau\in\mathbb{R}$ over the output of $f(RDR)$, e.g., a percentage $0\leq\tau_{p}\leq 1$. The controller’s privacy preference are satisfied when $f(RDR)\geq\tau$. In the above example, if the controller sets $\tau_{p}=0.9$, their privacy preference are satisfied when $\epsilon=0.1$ since $\frac{10}{11}\geq 0.9$.
@@ 4.2. The Find-$\epsilon$-from-RDR Algorithm
We design the Find-$\epsilon$-from-RDR algorithm (Algorithm 1) to operationalize the key idea presented above. The algorithm provides a general framework for finding a suitable $\epsilon$ that aligns with a controller’s specific privacy preference on the within-dataset individuals. Instead of prescribing a single, fixed RDR metric, our approach allows the controller to define their own preference function over RDRs, $f(RDR)$, and a corresponding threshold, $\tau$. The algorithm then finds the largest candidate $\epsilon$ (preferring higher utility) that satisfies the controller’s stated preference, $f(RDR)\geq\tau$.
Input : DP mechanism $\mathcal{M}$, dataset $x$, query $q$, preference function $f$, threshold $\tau$, set of candidate $\epsilon$ values $\mathcal{E}$, failure parameter $\delta$
@@
1ex
$x^{\prime}\leftarrow\textsc{project-query-attributes}(x,q)$
$PIS\leftarrow map()$
for $x_{i}\in x^{\prime}$ do
  if $x_{i}\notin PIS$ then
     $PIS[x_{i}]\leftarrow\textsc{compute-per-instance-sensitivity}(x^{\prime},x_{i},q,M)$
  end if
end for
/* Controllers may observe the computed per-instance sensitivity of each individual first, and set $f$ and $\tau$ based on their privacy preferences */
for $\epsilon\in\mathcal{E}$ do
 $RDR\leftarrow list()$
  for $x_{i}\in x^{\prime}$ do
     $RDR_{i}\leftarrow\textsc{compute-rdr}(PIS[x_{i}],M,q,\epsilon,\delta)$
     $RDR.add(RDR_{i})$
  end for
 /* Check whether $f(RDR)$ satisfy controller’s privacy preference */
  if $f(RDR)\geq\tau$ then
    /* Found suitable $\epsilon$. Compute DP output using chosen $\epsilon$ */
     $o\leftarrow\mathcal{M}(x^{\prime},q,\epsilon,\delta)$
     return $o$
  end if
end for
@@
Choosing preference function. The central argument of our work is that the controller, who has complete knowledge of individuals’ data and is responsible for protecting those within-dataset individuals, is the best-positioned agent to define what constitutes an acceptable risk distribution. By first computing and observing the within-dataset individuals’ per-instance sensitivities (Algorithm 1, lines 1 - 1, which will be used to compute individuals’ RDRs later), the controller gains a concrete understanding of the baseline disclosure risk of each individual. Armed with this information, they can formulate a preference $f,\tau$ that is not abstract, but rather grounded in the specific characteristics of their data. This reframes the problem from the difficult task of interpreting a global, worst case $\epsilon$ to the more intuitive one of expressing a direct policy on the relative risks they are willing to accept for the individuals they protect.
The flexibility of the preference function is a key strength of our framework, allowing controllers to express a wide variety of privacy preferences. As shown in our previous example, a controller may wish to ensure that no single individual is disproportionately more at risk than any other. This can be expressed using the preference function $f(RDR)=RDR_{min}/RDR_{max}\geq\tau_{p}$, where $\tau_{p}$ is a value close to 1 (e.g., 0.95). Alternatively, a controller might be less concerned with outliers and more with ensuring the overall risk distribution is tight. This goal can be captured by minimizing variance: $f(RDR)=-Var(RDR)\geq-\tau_{var}$ (equivalent to $Var(RDR)\leq\tau_{var}$), for a small variance threshold $\tau_{var}$. In addition, a controller may be particularly concerned about protecting vulnerable subgroups (e.g., minorities in a demographic dataset). They could define a preference that compares the median risk of a minority group to that of the majority group: $f(RDR)=RDR_{median\_majority}/RDR_{median\_minority}\geq\tau_{g}$, ensuring the risks between groups remain balanced. The choice of the preference function $f$ and threshold $\tau$ is not an arbitrary task but a deliberative, socio-technical process that depends on the controller’s privacy goals and their understanding of the data, now complemented with the additional information of the per-instance sensitivities and RDRs of within-dataset individuals.
Detailed Algorithm. Algorithm 1 shows the pseudo-code for finding $\epsilon$ from RDR and releasing DP output $o$. The input includes the set of candidate $\epsilon$ values to choose from; for example, the values could range from 10 to 0.01 in a fixed interval. The algorithm first projects relevant query attributes from the dataset (line 1) to avoid unnecessary recomputation of duplicate data (e.g., if the query only involves the age attribute, the algorithm only needs to compute the RDRs of hundreds of unique ages, regardless of the original size of the dataset). Then, for each unique within-dataset individual’s data $x_{i}$, the algorithm computes its per-instance sensitivity, which depends on the query and which DP mechanism will be used (line 1 - 1); the per-instance sensitivity will be used to compute the RDR of each within-dataset individual. At this point, the controller may observe each within-dataset individual’s per-instance sensitivity and sets the preference function $f$ and corresponding threshold $\tau$ based on their privacy preference over the within-dataset individuals.
Next, the algorithm iterates over the set of candidate $\epsilon$ values in descending order (line 1). For each $\epsilon$, the algorithm computes the RDR of each individual (line 1 - 1). If $f(RDR)\geq\tau$, then the algorithm has found a suitable $\epsilon$ that satisfies the controller’s privacy preference, so it computes a DP output using the chosen $\epsilon$ and returns the output (line 1 - 1). Note that the algorithm does not release the chosen $\epsilon$, since $\epsilon$ is chosen based on data-dependent RDRs and only the controller can access the RDRs and chosen $\epsilon$.
@@
Dataflows and leakage. It is critically important to understand what information is released to whom. Figure 3 shows the involved dataflows in a scenario where a controller uses the Find-$\epsilon$-from-RDR algorithm to find $\epsilon$ and release query output to the analyst. Individuals send their data to the controller; this dataflow remains unchanged from the standard central DP deployment shown in Figure 1. The analyst submits a query to the controller. The controller expresses their privacy preference in terms of $\tau_{p}$ (and not in terms of $\epsilon$) and runs the algorithm to release an output produced by a DP mechanism (Laplace Mechanism or Gaussian Mechanism in our work) with an $\epsilon$ selected by the algorithm. At this point, no information has been returned to the analyst yet. The RDRs of within-dataset individuals are never released to the analyst. The DP output is only released with the controller’s approval. The chosen $\epsilon$ is also not released since $\epsilon$ is chosen based on data-dependent RDRs, and releasing it directly jeopardizes the DP guarantee. In Section 5, we explain a modification of the algorithm that permits finding and releasing $\epsilon$ while satisfying DP.
@@
Claim: The Find-$\epsilon$-from-RDR algorithm satisfies $(\epsilon,\delta)$-DP. The algorithm only releases the output $o$ computed using an $(\epsilon,\delta)$-DP mechanism ($\delta=0$ when $\mathcal{M}$ is Laplace Mechanism), where $\epsilon$ is found by the algorithm that satisfies the controller’s privacy preference. If no suitable $\epsilon$ is found, the algorithm does not release anything; in other words, $\epsilon=0$. Therefore, the algorithm satisfies $(\epsilon,\delta)$-DP.
[…]
@@
5. Releasing Epsilon Privately and Bounding Privacy Loss Across Queries